OpenStudy (anonymous):

The Fun Guys game rental store charges an annual fee of $5 plus $5.50 per game rented. The Game Bank charges an annual fee of $14 plus $2.50 per game. For how many game rentals will the cost be the same at both stores?

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

\[5+5.5x=14+2.5x\] solve for \(x\)

OpenStudy (anonymous):

\[3x+5=14\\ 3x=9\\x=3\] is the rest of the work

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

whats the cost of it?

OpenStudy (anonymous):

nvm got it